Activity:

0.01 kg of a certain perfect gas occupies a volume of 0.003m3 at a pressure of 7 bar and a temperature of 131°C. Calculate the molar mass of the gas. When the gas is allowed to expand until the pressure is 1 bar the final volume is 0.02m3. Calculate the final temperature.



1
Expert's answer
2021-07-12T12:16:54-0400

"p_1V_1=mRT_1 \\

R=\frac{p_1V_1}{mT_1} \\

T_1=131+273=404 \;K \\

R= \frac{7 \times 10^5 \times 0.003}{0.01 \times 404} = 520 \; N \cdot m/kg \\

R=\frac{R_0}{M} \\

M = \frac{R_0}{R} \\

M=\frac{8314}{520}=16"

Molecular weight = 16

"p_2V_2=mRT_2 \\

T_2 = \frac{p_2V_2}{mR} \\

T_2= \frac{1 \times 10^5 \times 0.02}{0.01 \times 520} = 384.5 \;K"

Final temperature "= 384.5-273=111.5 \;°C"